OpenStudy (karabella98):

MEDAL AND FAN George does 18 J of work to lift a 1 kg box at a constant speed. If he drops it, how fast will the box be going when it hits the ground? A. 18 m/s B. 1.8 m/s C. 6 m/s D. 36 m/s

OpenStudy (noelgreco):

W = 1/2 mv^2

OpenStudy (abhisar):

Yes, the work done on the box has been stored as potential energy. Applying conservation of mechanical energy when the box is dropped this potential energy will get converted into kinetic energy. \(\sf \therefore 18=\frac{1}{2}\times m \times v^2\) Plug in the values and calculate v.

OpenStudy (anonymous):

work= energy=KE 18=1/2*m*V^2 V=sqrt(18*2/1)=sqrt(36)=6m/s